# Elliptic Virtual Structure Constants and Gromov-Witten Invariants for Complete Intersections in Weighted Projective Space

### v1

@@ 3 Setup of Our Computation
×$\qquad\times\prod_{p=1}^{m}\left(\frac{\prod_{j=1}^{d}e_{k_{p}}(z_{j-1},z_{j})}{\prod_{j=1}^{d-1}k_{p}z_{j}(2z_{j}-z_{j-1}-z_{j+1})}\right)\frac{1}{\prod_{j=1}^{d}q(z_{j-1},z_{j})}\left(\prod_{l=0}^{N-m-1}\left(\sum_{i=1}^{d}w_{l}(z_{i-1},z_{i})\right)^{n_{l}}\right).$ (3.16)
@@ 4.2 Residue Integrals for Elliptic Virtual Structure Constants
$\frac{1}{(w-z_{0})^{2}q(w,z_{0})(q(z_{0},z_{0}))^{f-1}}\left(\prod_{l=1}^{m}\frac{1}{(k_{l}z_{0})^{l-1}}\frac{e_{k_{l}}(w,z_{0})}{k_{l}w}\left(\frac{e_{k_{l}}(z_{0},z_{0})}{k_{l}z_{0}}\right)^{f-1}\right)$
×$\times\left(\prod_{i=1}^{l}\frac{\prod_{l=1}^{m}e_{k_{l}}(z_{0},z_{i,1})}{q(z_{0},z_{i,1})(z_{i,1}-z_{0})}\right)\left(\prod_{i=1}^{l}\left(\prod_{j=1}^{d_{i}-1}\frac{\prod_{l=1}^{m}e_{k_{l}}(z_{i,j},z_{i,j+1})}{q(z_{i,j},z_{i,j+1})(2z_{i,j}-z_{i,j-1}-z_{i,j+1})\left(\prod_{l=1}^{m}k_{l}z_{i,j}\right)}\right)\right)$
@@ 5 $99328$ $40960$
4 (0,6) $384$ $-\frac{57856}{3}$ 00 $-\frac{18667312}{3}$
4 (2,5) $2560$ $-\frac{57856}{3}$ 00 $-\frac{101879272}{3}$
4 (4,4) $18944$ $-\frac{57856}{3}$ $256$ $-\frac{555449168}{3}$
4 (6,3) $163840$ $-\frac{57856}{3}$ $3584$ $-\frac{3026251616}{3}$
4 (8,2) $1583104$ $-\frac{57856}{3}$ $43008$ $-\frac{16485590720}{3}$
4 (10,1) $16687104$ $-\frac{696320}{3}$ $512000$ $-\frac{89806527616}{3}$
4 (12,0) $189358080$ $-\frac{10141696}{3}$ $6246400$ $-163085218816$
@@
In this Appendix, we prove Proposition 1. The proof is done fundamentally in the same way as the corresponding proposition in [9], i.e., we have only to show that the residue of $f_{\Gamma}$ for type (iii) graphs at $w=z_{0}$ vanishes. If $a_{i}=1$, then $q(x,y)=\prod_{j=1}^{a_{i}-1}(jx+(a_{i}-j)y):=1$. Then we assume
@@
or $a_{1}=\cdots=a_{r}=1$. We also assume that $a_{i}\;\;(i=r+1,r+2,\cdots,N)$ is greater than $1$. From the Calabi-Yau condition, we have $r+\sum_{i=r+1}^{N}a_{i}=\sum_{j=1}^{m}k_{m}$.
$f_{\Gamma}=\frac{\text{sym}(\sigma)}{24(\prod_{i=1}^{r}a_{i})^{d}(z_{0})^{N(f-1)}}\left(\prod_{i=1}^{l}\left(\prod_{j=1}^{d_{i}}\frac{1}{(z_{i,j})^{N}}\right)\right)\left(-\frac{N-m}{N}\frac{1}{w^{N}}-\frac{N+m}{N}\frac{1}{(z_{0})^{N}}\right)$
×$\times\frac{1}{(w-z_{0})^{2}q(w,z_{0})(q(z_{0},z_{0}))^{f-1}}$
×$\times\left(\prod_{l=1}^{m}\frac{1}{(k_{l}z_{0})^{l-1}}\frac{e_{k_{l}}(w,z_{0})}{k_{l}w}\left(\frac{e_{k_{l}}(z_{0},z_{0})}{k_{l}z_{0}}\right)^{f-1}\right)\left(\prod_{i=1}^{l}\frac{\prod_{l=1}^{m}e_{k_{l}}(z_{0},z_{i,1})}{q(z_{0},z_{i,1})(z_{i,1}-z_{0})}\right)$
×$\times\left(\prod_{i=1}^{l}\left(\prod_{j=1}^{d_{i}-1}\frac{\prod_{l=1}^{m}e_{k_{l}}(z_{i,j},z_{i,j+1})}{q(z_{i,j},z_{i,j+1})(2z_{i,j}-z_{i,j-1}-z_{i,j+1})\left(\prod_{l=1}^{m}k_{l}z_{i,j}\right)}\right)\right)$
@@
Here, $g(z)$ in the last line represents the factor which does not contain $w$. We note here the following equalities:
@@
$\frac{d}{dw}q(w,z_{0})|_{w=z_{0}}=\sum_{i=r+1}^{N}\left(\prod_{\begin{subarray}{c}l=r+1\\ l\neq i\end{subarray}}^{N}(a_{l}z_{0})^{a_{l}-1}\right)(a_{i}z_{0})^{a_{i}-2}\frac{a_{i}(a_{i}-1)}{2}$ (5.135)
Using these, the residue of $f_{\Gamma}$ at $w=z_{0}$ is computed as follows.
$\lim_{w\to z_{0}}\frac{d}{dw}(f_{\Gamma})=\frac{d}{dw}\left[\left(-\frac{N-m}{N}\frac{1}{w^{N}}-\frac{N+m}{N}\frac{1}{(z_{0})^{N}}\right)\frac{1}{q(w,z_{0})}\left(\prod_{l=1}^{m}\frac{e_{k_{l}}(w,z_{0})}{k_{l}w}\right)\right]_{w=z_{0}}g(z)$
$=g(z)\left[\left(\frac{(N-m)}{(z_{0})^{N+1}}\right)\frac{1}{q(z_{0},z_{0})}\left(\prod_{l=1}^{m}\frac{e_{k_{l}}(z_{0},z_{0})}{k_{l}z_{0}}\right)+\frac{2}{(z_{0})^{N}}\left[\frac{\frac{d}{dw}q(w,z_{0})}{(q(w,z_{0}))^{2}}\right]_{w=z_{0}}\left(\prod_{l=1}^{m}\frac{e_{k_{l}}(z_{0},z_{0})}{k_{l}z_{0}}\right)\right.$
@@
$=g(z)\left[\frac{(N-m)}{(z_{0})^{N+1}}\frac{1}{\prod_{i=r+1}^{N}(a_{i}z_{0})^{a_{i}-1}}\left(\prod_{l=1}^{m}(k_{l}z_{0})^{k_{l}}\right)\right.$
@@
$\left.-\frac{2\prod_{l=1}^{m}(k_{l}z_{0})^{k_{l}}}{(z_{0})^{N}\prod_{i=1}^{r}(a_{i}z_{0})^{a_{i}-1}}\sum_{i=1}^{m}\frac{k_{i}-1}{2z_{0}}\right]$ (5.137)
$=g(z)\frac{\prod_{l=1}^{m}(k_{l}z_{0})^{k_{l}}}{(z_{0})^{N}\prod_{i=r+1}^{N}(a_{i}z_{0})^{a_{i}-1}}\left[\frac{N-m}{z_{0}}+\sum_{i=1}^{r}\frac{(a_{i}-1)}{z_{0}}-\sum_{i=1}^{m}\frac{k_{i}-1}{z_{0}}\right]$
$=g(z)\frac{\prod_{l=1}^{m}(k_{l}z_{0})^{k_{l}}}{(z_{0})^{N+1}\prod_{i=r+1}^{N}(a_{i}z_{0})^{a_{i}-1}}\left[N-m+\sum_{i=1}^{r}a_{i}-r-\sum_{i=1}^{m}k_{i}+m\right]=0.$ (5.138)
@@ Acknowledgment
Our research is partially supported by JSPS Grant No. 22K03289. Research of M.J. is also partially supported by JSPS Grant No. 24H00182.

### v3

@@ 3 Setup of Our Computation
×$\qquad\times\prod_{j=1}^{d}\left(\frac{\prod_{p=1}^{m}e_{k_{p}}(z_{j-1},z_{j})}{\left(\prod_{p=1}^{m}k_{p}z_{j}\right)(2z_{j}-z_{j-1}-z_{j+1})}\right)\left(\prod_{p=1}^{m}k_{p}z_{d}\right)\frac{1}{\prod_{j=1}^{d}q(z_{j-1},z_{j})}$
×$\qquad\times\left(\prod_{l=0}^{N-m-1}\left(\sum_{i=1}^{d}w_{l}(z_{i-1},z_{i})\right)^{n_{l}}\right).$ (3.16)
@@ 4.2 Residue Integrals for Elliptic Virtual Structure Constants
$\frac{1}{(w-z_{0})^{2}q(w,z_{0})(q(z_{0},z_{0}))^{f-1}}\left(\prod_{p=1}^{m}\frac{1}{(k_{p}z_{0})^{l-1}}\frac{e_{k_{p}}(w,z_{0})\left(e_{k_{p}}(z_{0},z_{0})\right)^{f-1}}{k_{p}w\left(k_{p}z_{0}\right)^{f}}\right)$
×$\times\left(\prod_{i=1}^{l}\frac{\prod_{p=1}^{m}e_{k_{p}}(z_{0},z_{i,1})}{q(z_{0},z_{i,1})(z_{i,1}-z_{0})}\right)\left(\prod_{i=1}^{l}\left(\prod_{j=1}^{d_{i}-1}\frac{\prod_{p=1}^{m}e_{k_{p}}(z_{i,j},z_{i,j+1})}{q(z_{i,j},z_{i,j+1})(2z_{i,j}-z_{i,j-1}-z_{i,j+1})\left(\prod_{p=1}^{m}k_{p}z_{i,j}\right)}\right)\right)$
@@ 5 $99328$ $40960$
4 (0,6) $384$ $-160$ 00 $-\frac{18667312}{3}$
4 (2,5) $2560$ $-\frac{3200}{3}$ 00 $-\frac{101879272}{3}$
4 (4,4) $18944$ $-\frac{22912}{3}$ $256$ $-\frac{555449168}{3}$
4 (6,3) $163840$ $-\frac{194048}{3}$ $3584$ $-\frac{3026251616}{3}$
4 (8,2) $1583104$ $-\frac{1849856}{3}$ $43008$ $-\frac{16485590720}{3}$
4 (10,1) $16687104$ $-6440960$ $512000$ $-\frac{89806527616}{3}$
4 (12,0) $189358080$ $-72652800$ $6246400$ $-163085218816$
@@
In this appendix, we prove Proposition 1. The proof fundamentally follows the same approach as the corresponding proposition in [9]; specifically, we only need to show that the residue of $f_{\Gamma}$ for type (iii) graphs at $w=z_{0}$ vanishes. If $a_{i}=1$, then $q(x,y)=\prod_{j=1}^{a_{i}-1}(jx+(a_{i}-j)y):=1$.
We assume
@@
which implies $a_{1}=\cdots=a_{r}=1$. We also assume that $a_{i}>1$ for $i\in\{r+1,r+2,\cdots,N\}$. From the Calabi-Yau condition, we have $r+\sum_{i=r+1}^{N}a_{i}=\sum_{j=1}^{m}k_{m}$.
$f_{\Gamma}$ $=\frac{\text{Sym}(\sigma)}{24(\prod_{i=1}^{N}a_{i})^{d}(z_{0})^{N(f-1)}}\left(\prod_{i=1}^{l}\left(\prod_{j=1}^{d_{i}}\frac{1}{(z_{i,j})^{N}}\right)\right)\left(-\frac{N-m}{N}\frac{1}{w^{N}}-\frac{N+m}{N}\frac{1}{(z_{0})^{N}}\right)$
$\frac{1}{(w-z_{0})^{2}q(w,z_{0})(q(z_{0},z_{0}))^{f-1}}\left(\prod_{p=1}^{m}\frac{1}{(k_{p}z_{0})^{l-1}}\frac{e_{k_{p}}(w,z_{0})\left(e_{k_{p}}(z_{0},z_{0})\right)^{f-1}}{k_{p}w\left(k_{p}z_{0}\right)^{f}}\right)$
×$\times\left(\prod_{i=1}^{l}\frac{\prod_{p=1}^{m}e_{k_{p}}(z_{0},z_{i,1})}{q(z_{0},z_{i,1})(z_{i,1}-z_{0})}\right)\left(\prod_{i=1}^{l}\left(\prod_{j=1}^{d_{i}-1}\frac{\prod_{p=1}^{m}e_{k_{p}}(z_{i,j},z_{i,j+1})}{q(z_{i,j},z_{i,j+1})(2z_{i,j}-z_{i,j-1}-z_{i,j+1})\left(\prod_{p=1}^{m}k_{p}z_{i,j}\right)}\right)\right)$
×$\times\prod_{a=2}^{N-m-1}\left(w_{a}(w,z_{0})+(f-1)w_{a}(z_{0},z_{0})+\sum_{i=1}^{l}\sum_{j=1}^{d_{i}}w_{a}(z_{i,j-1},z_{i,j})\right)^{n_{a}}.$
@@
Here, $g(z)$ in the last line represents the factor that does not contain $w$.
We note the following equalities:
@@
$\frac{d}{dw}q(w,z_{0})\Big{|}_{w=z_{0}}$ $=\sum_{i=r+1}^{N}\left(\prod_{\begin{subarray}{c}l=r+1\\ l\neq i\end{subarray}}^{N}(a_{l}z_{0})^{a_{l}-1}\right)(a_{i}z_{0})^{a_{i}-2}\frac{a_{i}(a_{i}-1)}{2}$ (5.135)
Using these equalities, the residue of $f_{\Gamma}$ at $w=z_{0}$ is computed as follows:
$\lim_{w\to z_{0}}\frac{d}{dw}(f_{\Gamma})$
$=\frac{d}{dw}\left[\left(-\frac{N-m}{N}\frac{1}{w^{N}}-\frac{N+m}{N}\frac{1}{(z_{0})^{N}}\right)\frac{1}{q(w,z_{0})}\left(\prod_{l=1}^{m}\frac{e_{k_{l}}(w,z_{0})}{k_{l}w}\right)\right]_{w=z_{0}}g(z)$
$=g(z)\left[\left(\frac{N-m}{(z_{0})^{N+1}}\right)\frac{1}{q(z_{0},z_{0})}\left(\prod_{l=1}^{m}\frac{e_{k_{l}}(z_{0},z_{0})}{k_{l}z_{0}}\right)+\frac{2}{(z_{0})^{N}}\left[\frac{\frac{d}{dw}q(w,z_{0})}{(q(w,z_{0}))^{2}}\right]_{w=z_{0}}\left(\prod_{l=1}^{m}\frac{e_{k_{l}}(z_{0},z_{0})}{k_{l}z_{0}}\right)\right.$
@@
$=g(z)\left[\frac{N-m}{(z_{0})^{N+1}}\frac{1}{\prod_{i=r+1}^{N}(a_{i}z_{0})^{a_{i}-1}}\left(\prod_{l=1}^{m}(k_{l}z_{0})^{k_{l}}\right)\right.$
@@
$\left.-\frac{2\prod_{l=1}^{m}(k_{l}z_{0})^{k_{l}}}{(z_{0})^{N}\prod_{i=r+1}^{N}(a_{i}z_{0})^{a_{i}-1}}\sum_{i=1}^{m}\frac{k_{i}-1}{2z_{0}}\right]$ (5.137)
$=g(z)\frac{\prod_{l=1}^{m}(k_{l}z_{0})^{k_{l}}}{(z_{0})^{N}\prod_{i=r+1}^{N}(a_{i}z_{0})^{a_{i}-1}}\left[\frac{N-m}{z_{0}}+\sum_{i=r+1}^{N}\frac{a_{i}-1}{z_{0}}-\sum_{i=1}^{m}\frac{k_{i}-1}{z_{0}}\right]$
$=g(z)\frac{\prod_{l=1}^{m}(k_{l}z_{0})^{k_{l}}}{(z_{0})^{N+1}\prod_{i=r+1}^{N}(a_{i}z_{0})^{a_{i}-1}}\left[N-m+\sum_{i=r+1}^{N}a_{i}-(N-r)-\sum_{i=1}^{m}k_{i}+m\right]=0.$ (5.138)
@@ Acknowledgment
This research is partially supported by JSPS Grant No. 22K03289. The research of M.J. is also partially supported by JSPS Grant No. 24H00182.